OpenStudy (anonymous):

graph the following 3x+4y=8

OpenStudy (anonymous):

solve for y. so y=2-3/4x when x=0, y=2 so (0,2) is one point and when y=0, x=8/3 since this is the equation of a line just connect the two known points, namely the x and y intercepts.

OpenStudy (anonymous):

oh totally. :P